OpenStudy (anonymous):

How many pieces of wire 2 1/2in. long can be cut from a piece 2ft. long?

jimthompson5910 (jim_thompson5910):

Hints: 2 1/2 inches = 2 + 1/2 2 1/2 inches = 4/2 + 1/2 2 1/2 inches = 5/2 inches 2 1/2 inches = 2.5 inches ------------- 2 ft = 2*12 inches 2 ft = 24 inches

OpenStudy (anonymous):

So, is the answer 24 inches?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

2 ft is the same as 24 inches

jimthompson5910 (jim_thompson5910):

2 1/2 inches = 2.5 inches

jimthompson5910 (jim_thompson5910):

to find the answer, you would divide the two

OpenStudy (anonymous):

Thanks